How many moles of sulfur atoms are present in a sample that contains 2.16 moles of disulfur decafluoride, S2F10 ?


...................mol S


How many moles of S2F10  molecules are present in a sample that contains 3.07 moles of sulfur atoms?


.............mol S2F10


Please with steps.


1
Expert's answer
2021-02-28T06:20:37-0500
  1. n(S)= 2*2.16=4.32 mol
  2. n(S2F10)=3.07/2= 1.535 mol